OpenStudy (anonymous):

will give medal and fan whats x+3y=15

pooja195 (pooja195):

wee i figure since your trying to find the Y value subtract x so that your equation if 3y=X-15 and then divide by 3

pooja195 (pooja195):

x + 3y = 15 3y = 15- x y = 5 - 1/3 x

pooja195 (pooja195):

Do you get it?

OpenStudy (anonymous):

yes